OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. x2 + 8x - 5 = 0

OpenStudy (anonymous):

- b +/- root(b^2 - 4ac) x = ---------------------- 2a - (8) +/- root(8^2 - 4(1)(-5) -8 +/- root(64 + 20) x = -------------------------- = -------------------- 2*1 2 -8 +/- root84 -8 +/- 2root21 x = -------------- = -------------- 2 2

OpenStudy (anonymous):

-8 + 2root21 -8 - 2root21 x = ------------- or ------------ 2 2 2(-4 + root21) 2(-4 - root21) x = --------------- or ------------- 2 2

OpenStudy (anonymous):

so final answer \[x = -4 + \sqrt{21}\] or

OpenStudy (anonymous):

\[x = -4 -\sqrt{21}\]